A couple invests $8,705.00 today in a money market fund that pays 6.00% per year. What is the value of this account in exactly one year?

A loan shark offers to give you $5,274.00 today, but, in exchange, he wants you to pay him $6,876.00 in one year. What is the rate of interest being charged on the account?

A young graduate looks to save money to buy a house 4.00 years from today. He is somewhat conservative and will invest his money in a bond fund that pays 4.00% annually. The graduate invests $10,886.00 today. How much will his account be worth in 4.00 years?

A couple is moving to San Diego, California in 4.00 years when they finish medical school. They anticipate needing $100,000.00 to make a down payment on a condo in the area they want to work. The couple has a money market account that earns 4.00% per year. How much should they set aside today to make the future payment?
Answer #1

The solution to the given question is provided below for your reference

1.
=8705*1.06=9227.3

2.
=6876/5274-1=30.3754%

3.
=10886*1.04^4=12735.08028

4.
=100000/1.04^4=85480.4191